# A Fuzzy-based Approach to Predict Human Interaction by Functional Near-Infrared Spectroscopy

### v1

@@ Abstract
The paper introduces a Fuzzy-based Attention (Fuzzy Attention Layer) mechanism, a novel computational approach to enhance the interpretability and efficacy of neural models in psychological research. The proposed Fuzzy Attention Layer mechanism is integrated as a neural network layer within the Transformer Encoder model to facilitate the analysis of complex psychological phenomena through neural signals, such as those captured by functional Near-Infrared Spectroscopy (fNIRS). By leveraging fuzzy logic, the Fuzzy Attention Layer is capable of learning and identifying interpretable patterns of neural activity. This capability addresses a significant challenge when using Transformer: the lack of transparency in determining which specific brain activities most contribute to particular predictions. Our experimental results demonstrated on fNIRS data from subjects engaged in social interactions involving handholding reveal that the Fuzzy Attention Layer not only learns interpretable patterns of neural activity but also enhances model performance. Additionally, the learned patterns provide deeper insights into the neural correlates of interpersonal touch and emotional exchange. The application of our model shows promising potential in deciphering the subtle complexities of human social behaviors, thereby contributing significantly to the fields of social neuroscience and psychological AI.
@@ I Introduction
Figure 1: Overall illustration of decoding handholding by hyperscanning fNIRS signals. A: A demonstration sample showing paired fNIRS signals. The brain image displays the prefrontal cortex’s maximum concentration of HbO and HbR. The second plot illustrates how the average signal across channels for HbO and HbR changes over time. B: The main structure of our proposed model. C: The center, which is identified by the Fuzzy Attention Layer. D: The structure of the Fuzzy Attention Layer. E: The firing strength of the demonstration sample.
@@
This paper introduces an interdisciplinary, bottom-up approach with an explainable fuzzy logic-based model to support psychological and neuroscientific theories. We propose the Fuzzy Rule-Based Attention (Fuzzy Attention Layer), illustrated in Figure 1D, which patterns as fuzzy sets and form fuzzy rules for decision-making, which confers the neural network with human-interpretable structures[28]. Designed for data-sparse environments like functional Near-Infrared Spectroscopy (fNIRS), it automatically learns the neural activity patterns in a human-interpretable manner, providing evidence for human interaction behavior and enhancing performance compared to traditional dot-product attention in neural datasets.
@@
    Proposing Fuzzy Attention Layer: A fuzzy logic-based attention mechanism to improve the interpretability of a Transformer without sacrificing its performance.
@@
    Revealing underlying Neural Patterns: By examining the fuzzy rules learned by the Fuzzy Attention Layer, we uncover the underlying neural patterns of specific human-to-human interactions.
Figure 2: A: Channel-first structure: The T-values illustrate the differences between handholding (HH) and non-handholding (NHH) conditions across 10 different rules. B: Time-first structure: The T-values illustrate the differences between HH and NHH conditions across 10 rules. C: The #5 centre. D: The placement of fNIRS sensors (sources and detectors) and channels. E: Statistical analysis results showing the (de-)similarity between the latent variables of two individuals. ***:p<0.001
@@ II-B Fuzzy Inference Systems
Rooted in Fuzzy Logic, Fuzzy Inference Systems (FISs), particularly the Takagi–Sugeno–Kang (TSK) inference system, have collaborated with state-of-the-art AI technology and yielded significant advances [28]. FISs are composed of several IF-THEN rules, as follows:
@@
In this study, we take a further step by combining attention calculation with FIS, thus proposing the Fuzzy Attention Layer. This layer integrates FIS’s internal degrees of freedom into the Transformer architecture. It provides better approximation ability for brain signal data and higher transparency for interpretability analysis. We will discuss the Fuzzy Attention Layer in the next section.
@@ III Methodology
The overall process of the Fuzzy Attention Layer is illustrated in Figure 1B, where the fNIRS data of two individuals in one pair are put into the same feature extractor and extract the embedding of the neural responses in the same embedding space.
@@ III-A Task Definition
Given the fNIRS data recorded from two individuals observing the same image under two distinct conditions—handholding and not handholding—our task is to develop a classifier capable of discerning the participants’ interaction state. The fNIRS data, denoted as $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ for each participant, are obtained while simultaneously engaging with the visual stimulus. The classifier aims to predict the binary label $H\in\{0,1\}$, where $H=1$ corresponds to a handholding condition, and $H=0$ indicates no handholding. The challenge lies in capturing the subtle physiological synchronies and disparities influenced by the interpersonal touch encoded in the hemodynamic responses as measured by fNIRS. This task extends upon the emerging discourse in social neuroscience regarding the neural underpinnings of human physical interactions as explored in Section II-A. Our approach leverages a paired-sample framework, incorporating both within- and between-subject variations to accurately classify the interaction states, thus advancing the understanding of neural connectivity patterns associated with social touch.
@@ III-B1 Proposition 1: Self-Attention as the Cosine Component of a Fuzzy Set
$=\rm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{(x_{i,d}-m_{r,d})^{2}}{2\sigma_{r,d}^{2}}\right)$
$=\rm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{x_{i,d}^{2}+m_{r,d}^{2}-2x_{i,d}m_{r,d}}{2\sigma_{r,d}^{2}}\right)$
$=\rm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{x_{i,d}^{2}+m_{r,d}^{2}}{2\sigma_{r,d}^{2}}+\sum_{d=1}^{D}\frac{x_{i,d}m_{r,d}}{\sigma_{r,d}^{2}}\right)$ (4)
where $r$ is the index of the fuzzy rule among the total $R$ rules, $x_{d}$ is the input at dimension $d$, $m_{r,d}$ and $\sigma_{r,d}$ represent the centers and widths of the fuzzy rule at dimension $d$. Similarly, the formula for attention weights can be written as:
@@
$=\text{softmax}_{i,j}\left(\sum_{d=1}^{D}\frac{\mathbf{x}_{i,d}\mathbf{k}_{j,d}}{\sqrt{d_{k}}}\right).$ (5)
@@
Therefore, the Fuzzy Attention Layer enables flexible modeling of input patterns, allowing for analysis and learning of key information, and includes the mode as a contribution to attention.
@@
Fig. 3 provides an intuitive illustration of the differences between fuzzy attention(A) and dot-product attention scores(B) based on 2-dimensional inputs. The graphs reveal distinct characteristics of these attention mechanisms. In fuzzy attention, the surfaces exhibit convex shapes, indicating the presence of a maximum scored point. This point offers valuable insights into the patterns learned by the attention model. Conversely, dot-product attention scores form flat planes. When the input query aligns with the key, the attention scales infinitely with the query value. While this orientation-based scoring may be effective in NLP tasks, it proves inefficient for sparse data patterns. Putting simply, cosine similarity yields high scores when the correct orientation is identified and low scores for other orientations. In contrast, the fuzzy distance matrix exhibits high values in certain neighborhoods and low values in remote areas.
@@ III-C Fuzzy Transformer for fNIRS modeling
Our model comprises two main components: the feature extractor and the classification head. The feature extractor follows the structure of a Transformer Encoder, where the attention layers are replaced by a Fuzzy Attention Layer. As depicted in Figure 1B, this extractor is utilized for both participants in a pair. The output embedding from the encoder is concatenated and then processed by the classification head.
@@
where the $Q(x)$ s the linear projection from the input sequence, and $m$ and $\sigma$ are the Gaussian membership parameters, and $u$ the consequent of the fuzzy rule. From the attention view, the fuzzy attention score replaces the attention score matrix from the dot product similarity to an L2 distance weighted by $\sigma$. This attention score has been observed to be beneficial in fNIRS data modeling. The objective is to classify whether individuals are holding hands. We employ the cross-entropy loss, which is defined as
$\text{loss}(y_{o,c},p_{o,c})=-\sum_{c=1}^{M}y_{o,c}\log(p_{o,c})$ (13)
where $y_{o,c}$ is the true label, $p_{o,c}$ is the predicted probability for class $c$, and $M$ represents the total number of classes in the classification problem.
@@ IV Experiments and Results
TABLE I: Performance comparison of Fuzzy Attention Layer-based transformer against LSTM and vanilla transformer. The evaluation is conducted on two datasets with several matrices listed.
@@
Recognition channel-first Fuzzy 67.42% 69.39% 68.92% 69.15% 67.31% 63.94%
LSTM 60.39% 66.69% 61.39% 63.93% 60.04% 58.47%
Transformer 59.24% 53.92% 63.23% 58.20% 59.54% 58.35%
time-first Fuzzy 77.77% 82.30% 77.03% 79.58% 77.52% 72.71%
LSTM 62.24% 69.04% 62.87% 65.81% 61.87% 59.70%
Transformer 76.58% 75.55% 79.03% 77.25% 76.64% 72.57%
Rating channel-first Fuzzy 65.26% 61.07% 70.34% 65.38% 65.60% 63.86%
LSTM 64.58% 65.73% 67.48% 66.59% 64.49% 62.76%
Transformer 62.76% 56.64% 68.55% 62.03% 63.24% 62.11%
time-first Fuzzy 79.32% 83.25% 79.29% 81.22% 79.01% 75.00%
LSTM 74.16% 74.24% 76.86% 75.53% 74.15% 70.89%
Transformer 78.01% 81.12% 78.62% 79.85% 77.76% 73.92%
@@ IV-A Dataset
To evaluate our proposed Fuzzy Attention Layer aimed at decoding brain signals, we implemented two datasets in our study, each with different experimental setups named Picture Recognition and Picture Rating. Both experiments involved pairs of participants who performed tasks under two conditions: with and without handholding. The experiments were structured to minimize interaction by seating participants back-to-back. Figure 2D illustrates the fNIRS channel locations.
@@ IV-A1 Picture Recognition
The study adopted a two-factor mixed experimental design focusing primarily on the variables of hand-holding (hand-holding vs. non-hand-holding) and image type (negative vs. neutral). In this experiment, participants arrived at the laboratory and were briefed about the study’s procedures without revealing the hypotheses. They were reassured about the safety of the experiment and provided written informed consent. Before the start of the experiment, each participant was isolated in separate rooms to fill out initial questionnaires. Preparation included donning necessary devices. The protocol consisted of three sessions of approximately 5 minutes each for relaxation and briefing, interspersed with two task sessions lasting around 20 minutes each. During the relaxation phases, participants were instructed to remain still, close their eyes, and relax. Task sessions involved emotional image recognition tasks performed both in hand-holding and non-hand-holding condition, with participants responding to emotional stimuli presented on the screens using a two-button keypad—pressing the left button for negative images and the right for neutral ones. Each image was displayed following a 2-second fixation cross, and the response period was followed by feedback lasting 1 second. Each trial ensured a minimum duration of 11 seconds, extending up to 13 seconds to facilitate relaxation.
@@ IV-A2 Picture Rating
Similarly, The study also adopted a two-factor mixed experimental design focusing primarily on the variables of hand-holding (hand-holding vs. non-hand-holding) and image type (negative vs. neutral). The design was organized into a Block design. The protocol was divided into two main blocks based on hand-holding conditions and 16 sub-blocks based on the type of image presented, with each sub-block containing 10 trials of either negative or neutral images. To control for sequence effects, the order of the image type blocks was systematically varied using an ABBA-style arrangement. Each trial commenced with a 2-second presentation of a fixation point, followed by a 2-second solitary display of the image. Participants had 2 seconds to rate the emotional valence (from 1, negative, to 9, positive) and arousal (from 1, calm, to 9, excited). Upon completing ratings, the next trial began. Breaks of 30 minutes and 5 minutes were observed after each image type and hand-holding block, respectively. Participants started with a 5-minute baseline data collection session to measure resting neural activity. Following non-hand-holding tasks, the experiment introduced the hand-holding condition, where participants held hands and repeated the image evaluation task. The study concluded with another 5-minute resting state data collection session to assess any neural changes post-experiment.
IV-A3 fNIRS Setup and Preprocessing

### v2

# A Fuzzy Logic-based Approach to Predict Human Interaction by Functional Near-Infrared Spectroscopy

@@ Abstract
This paper introduces the Fuzzy logic-based Attention (Fuzzy Attention Layer) mechanism, a novel computational approach designed to enhance the interpretability and efficacy of neural models in psychological research. The Fuzzy Attention Layer integrated into the Transformer Encoder model to analyze complex psychological phenomena from neural signals captured by functional Near-Infrared Spectroscopy (fNIRS). By leveraging fuzzy logic, the Fuzzy Attention Layer learns and identifies interpretable patterns of neural activity. This addresses a significant challenge in using Transformers: the lack of transparency in determining which specific brain activities most contribute to particular predictions. Our experimental results, obtained from fNIRS data engaged in social interactions involving handholding, reveal that the Fuzzy Attention Layer not only learns interpretable patterns of neural activity but also enhances model performance. Additionally, these patterns provide deeper insights into the neural correlates of interpersonal touch and emotional exchange. The application of our model shows promising potential in understanding the complex aspects of human social behavior, verify psychological theory with machine learning algorithms, thereby contributing significantly to the fields of social neuroscience and AI.
@@ I Introduction
Figure 1: Overall illustration of decoding handholding by hyperscanning fNIRS signals. A: A demonstration sample showing paired fNIRS signals. The brain image displays the prefrontal cortex’s maximum concentration of HbO and HbR. The second plot illustrates how the average signal across channels for HbO and HbR changes over time. B: The main structure of our proposed model. Add&Norm: A residual connection followed by layer normalization. C: The center, which is identified by the Fuzzy Attention Layer. D: The structure of the Fuzzy Attention Layer. E: The firing strength of the demonstration sample.
@@
This paper introduces an interdisciplinary, bottom-up approach with an explainable fuzzy logic-based model to support psychological and neuroscientific theories. We propose the Fuzzy logic-Based Attention (Fuzzy Attention Layer), illustrated in Fig. 1D.
Designed for data-sparse environments like functional Near-Infrared Spectroscopy (fNIRS) [28, 29], it automatically learns the neural activity patterns in a human-interpretable manner, providing evidence for human interaction behavior and enhancing performance compared to traditional dot-product attention in neural datasets.
@@
This article is structured as follows. In Section II, we review recent work on brain decoding in human interactions and fuzzy inference systems. Section III describes the design and implementation of the proposed fuzzy attention layer and the fuzzy Transformer model. Experimental results and ablation studies are presented in Section IV to demonstrate the performance of the proposed approach and provide insights into the model’s behavior. Finally, conclusions and potential future research directions are discussed in Section V.
@@
    Proposing Fuzzy Attention Layer: The first to design an attention kernel based on fuzzy theory to improve the interpretability of a Transformer without sacrificing its performance.
@@
    Revealing underlying Neural Patterns: By examining the fuzzy rules learned by the Fuzzy Attention Layer, we uncover the underlying neural patterns of specific human-to-human interactions , and it connects machine learning algorithms with psychological theories.
Figure 2: A: The T-values, which illustrate the differences between handholding (HH) and non-handholding (NHH) conditions across 10 different rules, are compared within the channel-first structure. B: The T-values, which illustrate the differences between HH and NHH conditions across 10 rules, are compared within the time-first structure. C: The centre prototype for rule #5. D: The placement of fNIRS sensors (sources and detectors) and channels. E: Statistical analysis results showing the (de-)similarity between the latent variables of two individuals. ***:p<0.001
@@ II-B Fuzzy Inference Systems
This paper design the FIS based on the Takagi–Sugeno–Kang (TSK) inference system, which composed of several IF-THEN rules, as follows:
@@
In this study, we take a further step by combining attention calculation with FIS, thus proposing the Fuzzy Attention Layer. Details of this implementation is discussed in the next section.
@@ III-A Task Definition
To explore the participants’ interaction state, the experiment asked two subjects wearing fNIRS caps to watch the same image together under two conditions as the binary label $H\in\{0,1\}$, where $H=1$ corresponds to a handholding condition, and $H=0$ indicates not-handholding. The fNIRS data, denoted as $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ for each participant, are obtained while simultaneously engaging with the visual stimulus. The classifier aims to predict the label and capture the subtle physiological synchronization and disparities influenced by the interpersonal touch encoded in the hemodynamic responses as measured by fNIRS. This task extends upon the emerging discourse in social neuroscience regarding the neural underpinnings of human physical interactions as explored in Section II-A. Our approach leverages a paired-sample framework to enhance the classification of interaction states by incorporating both within- and between-subject variations. This method advances our understanding of the neural connectivity patterns associated with social touch. Specifically, within-subject variation involves extracting patterns using the model’s encoder, which processes each subject’s data separately. These patterns are then combined with those from paired subjects to analyze between-subject variations, ultimately making the final decision.
@@ III-B1 Proposition 1: Self-Attention as the Cosine Component of a Fuzzy Set
$=\mathrm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{(x_{i,d}-m_{r,d})^{2}}{2\sigma_{r,d}^{2}}\right)$
$=\mathrm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{x_{i,d}^{2}+m_{r,d}^{2}-2x_{i,d}m_{r,d}}{2\sigma_{r,d}^{2}}\right)$
$=\mathrm{softmax}_{i,r}\left(-\sum_{d=1}^{D}\frac{x_{i,d}^{2}+m_{r,d}^{2}}{2\sigma_{r,d}^{2}}+\sum_{d=1}^{D}\frac{x_{i,d}m_{r,d}}{\sigma_{r,d}^{2}}\right)$ (4)
where $x_{d}$ is the input at dimension $d$, $m_{r,d}$ and $m_{r,d}$ represent the centers and widths of the fuzzy rule at dimension $d$.The parameters $m_{r,d}$ and $\sigma_{r,d}$ are learnable and will evolve as training proceeds. Similarly to self-attention, the formula for the fuzzy attention weights can be written as:
@@
$=\text{softmax}_{i,j}\left(\sum_{d=1}^{D}\frac{{x}_{i,d}{k}_{j,d}}{\sqrt{d_{k}}}\right).$ (5)
@@
With such design, the flexibility of the length of the keys in Fuzzy Attention Layer enables the modeling of input patterns, allowing for analysis and learning of key information, and includes the mode as a contribution to attention. Fig. 3 provides an intuitive illustration of the differences between fuzzy attention(A) and dot-product attention scores(B) based on 2-dimensional inputs. The graphs reveal distinct characteristics of these attention mechanisms. In fuzzy attention, the surfaces exhibit convex shapes, indicating the presence of a maximum scored point. This point offers valuable insights into the patterns learned by the attention model. Conversely, dot-product attention scores form flat planes. When the input query aligns with the key, the attention scales infinitely with the query value. While this orientation-based scoring may be effective in NLP tasks, it proves inefficient for sparse data patterns. Putting simply, cosine similarity yields high scores when the correct orientation is identified and low scores for other orientations. In contrast, the fuzzy distance matrix exhibits high values in certain neighborhoods and low values in remote areas.
@@ III-C Fuzzy Transformer for fNIRS modeling
Our model comprises two main components: the feature extractor and the classification head. The feature extractor follows the structure of a Transformer Encoder, where the attention layers are replaced by a Fuzzy Attention Layer. As depicted in Fig. 1B, this extractor is utilized for both participants in a pair. The output embedding from the encoder is averaged in the second dimension $S$ and concatenated in the final dimension $C$. The shape of the concatenated vector is $2C$. Then, this concatenated vector is processed by the classification head, and the purpose of this concatenation is to combine information across multiple channels and from two subjects and make the final decision. The classification head is a 2-layer Multi-Layer Perception (MLP) model with a ReLU (Rectified Linear Unit) activation function. The output of the classification head consists of two nodes. This integration helps the model to capture and leverage the diverse signal patterns, enhancing its ability to differentiate between two conditions, thereby improving the overall performance and accuracy. The operations of the other basic layers in our encoder follow the standard Transformer architecture, with details available in [33].
@@
where the $Q(x)$ s the linear projection from the input sequence, and $m$ and $\sigma$ are the Gaussian membership parameters, and $u$ the consequent of the fuzzy rule. From the attention view, the fuzzy attention score replaces the attention score matrix from the dot product similarity to an L2 distance weighted by $\sigma$. This attention score has been observed to be beneficial in fNIRS data modeling.In this paper, the fuzzy rule count is set to 10, based on the results of the ablation study discussed in Section IV-E2.
This paper defines the objective function as a classification task to determine whether individuals are holding hands. We use cross-entropy loss, which is defined as:
$\text{loss}(y_{o,c},p_{o,c})=-\sum_{c=1}^{M}y_{o,c}\log(p_{o,c}),$ (13)
where $y_{o,c}$ represents the true label, $p_{o,c}$ is the predicted probability for class $c$, and $M$ is the total number of classes in the classification problem.
@@ IV Experiments and Results
TABLE I: Performance comparison of Fuzzy Attention Layer-based transformer against LSTM and vanilla transformer. The evaluation is conducted on two datasets with several matrices listed. * indicate significant differences between our proposed model and others, as assessed by t-tests (*$p<0.05$,**$p<0.01$, ***$p<0.001$$p<0.1$)
@@
Recognition channel-first Fuzzy 67.71% (0.96%) 70.50% (4.15%) 68.98% (1.78%) 69.64% (1.49%) 67.56% (0.99%) 64.11% (0.90%)
Transformer 60.55% (1.29%) *** 51.20% (5.83%) *** 67.71% (2.36%) 58.08% (3.47%) *** 61.30% (1.17%) *** 60.80% (0.87%) ***
LSTM 55.30% (0.98%) *** 59.35% (7.72%)* 58.97% (1.59%) *** 58.90% (3.04%) *** 54.92% (1.26%) *** 57.03% (0.75%) ***
time-first Fuzzy 79.85% (2.75%) 82.74% (2.37%) 79.83% (3.43%) 81.23% (2.34%) 79.68% (2.84%) 75.15% (3.08%)
Transformer 76.22% (1.20%)* 77.59% (3.21%)* 78.03% (0.60%) 77.78% (1.58%)* 76.11% (1.06%)* 72.57% (0.80%)
LSTM 65.47% (2.79%) *** 70.06% (3.29%) *** 64.02% (1.89%) *** 66.81% (0.89%) *** 62.32% (2.59%) *** 58.67% (4.14%) ***
Rating channel-first Fuzzy 66.98% (0.21%) 68.14% (3.48%) 69.76% (1.22%) 68.88% (1.16%) 66.89% (0.23%) 64.62% (0.32%)
Transformer 57.36% (1.75%) *** 57.49% (4.77%)** 60.04% (2.42%) *** 58.60% (2.09%) *** 57.35% (1.88%) *** 56.85% (1.22%) ***
LSTM 64.03% (0.93%) *** 69.65% (3.91%) 65.62% (1.56%)** 67.50% (1.34%) 63.58% (1.07%) *** 61.97% (0.82%) ***
time-first Fuzzy 78.01% (0.70%) 80.49% (2.41%) 79.00% (0.92%) 79.71% (0.94%) 77.81% (0.63%) 74.05% (0.56%)
Transformer 74.21% (3.96%). 78.78% (4.30%) 74.21% (4.97%). 76.31% (3.15%)* 73.96% (4.11%). 69.62% (3.89%)*
LSTM 74.56% (4.03%). 76.36% (2.21%)* 75.02% (2.59%)* 76.57% (2.42%)* 72.82% (4.46%)* 70.04% (2.24%)**
@@ IV-A Dataset
To evaluate our proposed Fuzzy Attention Layer aimed at decoding brain signals, we implemented two datasets in our study, each with different experimental setups named Picture Recognition and Picture Rating. Both experiments involved pairs of participants who performed tasks under two conditions: with and without handholding. The experiments were structured to minimize interaction by seating participants back-to-back. Fig. 2D illustrates the fNIRS channel locations. Both two datasets have been approved by the Ethical Review of Psychological Research in Henan Province Key Laboratory of Psychology and Behavior(20200702002).
Figure 4: The stimulus diagram of the experiment for two datasets. A: Picture Recognition Dataset and B: Picture Rating Dataset. This figure illustrates how data were collected under the handholding (HH) and non-handholding (NHH) conditions.
@@ IV-A1 Picture Recognition
The study adopted a three-factor mixed experimental design focusing primarily on the variables of relationships(friends vs. strangers), hand-holding (hand-holding vs. non-hand-holding), and image type (negative vs. neutral). In this experiment, participants arrived at the laboratory and were briefed about the study’s procedures without revealing the hypotheses. Before the start of the experiment, each participant was isolated in separate rooms to fill out initial questionnaires. Preparation included donning necessary devices. The protocol consisted of three sessions of approximately 5 minutes each for relaxation and briefing, interspersed with two task sessions lasting around 20 minutes each. During the relaxation phases, participants were instructed to remain still, close their eyes, and relax. Task sessions involved emotional image recognition tasks performed both in hand-holding and non-hand-holding conditions, with participants responding to emotional stimuli presented on the screens using a two-button keypad—pressing the left button for negative images and the right for neutral ones. Each image was displayed following a 2-second fixation cross, and the response period was followed by feedback lasting 1 second. Each trial ensured a minimum duration of 11 seconds, extending up to 13 seconds to facilitate relaxation. There are 40 events total for each affective image type in each handholding condition. The diagram of the experimental stimulus is presented in Fig. 4A. One example of the collected fNIRS data is shown in Fig. 1A.
@@ IV-A2 Picture Rating
Similarly, The study also adopted a three-factor mixed experimental design focusing primarily on the variables of relationships(friends vs. strangers), hand-holding (hand-holding vs. non-hand-holding), and image type (negative vs. neutral). The design was organized into a Block design. The protocol was divided into two main blocks based on hand-holding conditions and 16 sub-blocks based on the type of image presented, with each sub-block containing 10 trials of either negative or neutral images. To control for sequence effects, the order of the image type blocks was systematically varied using an ABBA-style arrangement. Each trial commenced with a 2-second presentation of a fixation point, followed by a 2-second solitary display of the image. Participants had 2 seconds to rate the emotional valence (from 1, negative, to 9, positive) and arousal (from 1, calm, to 9, excited). Upon completing ratings, the next trial began. Breaks of 30 minutes and 5 minutes were observed after each image type and hand-holding block, respectively. Participants started with a 5-minute baseline data collection session to measure resting neural activity. Following non-hand-holding tasks, the experiment introduced the hand-holding condition, where participants held hands and repeated the image evaluation task. The study concluded with another 5-minute resting state data collection session to assess any neural changes post-experiment. There are 40 events total for each affective image type in each handholding condition. The diagram of the experimental stimulus is presented in Fig. 4B.
IV-A3 Participants
Picture recognition includes 23 dyads of friends (Mean age = 20.156, SD = 2.022) and 26 dyads of strangers (Mean age =19.470, SD = 2.212), and the picture rating task includes 24 dyads of friends (Mean age = 19.786, SD = 1.704) and 23 dyads of strangers (Mean age = 19.932, SD = 1.404). All subjects are female and recruited through an online registration system. Based on their answer, our potential participants in the friend group had to have known each other for more than six months and consider each other one of their five best friends. Participants in the stranger group were paired randomly to ensure that they had never met before. All participants in this experiment were right-handed, heterosexual, had no history of mental illness, and had normal or corrected vision. Participants were asked to avoid staying up late and drinking alcohol the day before the experiment. They were reassured about the safety of the experiment and provided written informed consent. Participants will receive a compensation of 50 yuan approved by the Ethics Committee of the Henan Provincial Key Laboratory of Psychology and Behavior.
IV-A4 fNIRS Setup and Preprocessing